### v2

# Extremum Seeking with High-Order Lie Bracket Approximations: Achieving Exponential Decay Rate

Abstract
This paper focuses on the further development of the Lie bracket approximation approach for extremum seeking systems. Classical results in this area provide extremum seeking algorithms with exponential convergence rates for quadratic-like cost functions, and polynomial decay rates for cost functions of higher degrees. This paper proposes a novel control design approach that ensures the motion of the extremum seeking system along directions associated with higher-order Lie brackets, thereby ensuring exponential convergence for cost functions that are polynomial-like but with degree greater than two.
I Introduction
Consider the following control-affine system:
$\dot{x}=f_{0}(x)+\sum\limits_{i=1}^{m}\frac{1}{\varepsilon^{p_{i}}}f_{i}(x)u_{i}\left(k_{i}\frac{t}{\varepsilon}\right),\$ (1)
where $x\in\mathbb{R}^{n}$ is the state space vector, $p_{i}\in(0,1)$, $0<\varepsilon\ll 1$, $f_{0}$ is the drift (uncontrolled) vector field of the system, $f_{i}$ are the control vector fields, $u_{i}$ are the control inputs, $m\in Z^{+}$ is the number of control inputs, and $k_{i}\in Q_{>0}$ is a positive rational number. Control-affine systems of the form (1) characterize many real-world systems and applications, such as but not limited to robotic, multi-agent and flight dynamic systems (e.g., [1, 2, 3, 4, 5]). The geometric nature of systems of the form (1) allows the application of geometric control methods and techniques [6, 7] that are based on Lie-bracket-based approximations and analysis [8, 9, 10, 11, 12, 13, 14, 15]. A Lie bracket variation between the vector fields $f_{i}\text{ and }f_{j}$ is defined as:
$[f_{i},f_{j}]:=\frac{\partial f_{j}}{\partial x}f_{i}-\frac{\partial f_{i}}{\partial x}f_{j}.$ (2)
Lie-bracket-based approximations and analysis have been used to study motion planning of underactuated systems, including with nonholonomic constraints (e.g., [6, 10, 9, 16, 17, 18, 19, 20]) and model-free optimization and control via extremum seeking (ES) approaches [11, 12, 13, 14, 15]. In this paper, we focus on Lie bracket approximations relevant to ES systems.
Extremum seeking (ES) systems [21] are model-free and real-time dynamic optimization and control techniques that aim at stabilizing a dynamical system about the extremum point of an objective function, for which we have access to its measurements but not its expression; see [22] for a comprehensive review. ES methods referred to as classic-based (e.g., [23, 24, 25]) rely on classical averaging methods [26, 27] for analysis and design. On the other hand, control-affine ES methods [11, 12, 13, 14, 15] (the focus of this paper) rely on Lie-bracket approximations for analysis and design.
Now we briefly summarize the development of Lie bracket approximations of control-affine ES systems. For a particular case of (1) when $p_{i}=0.5,k_{i}=1\text{ for all }i$, a first-order Lie bracket system (LBS) approximation was given in [11], which also admits drift vector field and is defined not only for underactuated systems. In [28], the authors utilized the first-order LBS in [11] and provided a universal semi-global stabilizing control law based on Lyapunov function candidates; they also [29] extended the first-order LBS in [11] to a non-smooth setting. Different control law from [11] with bounded update rate was provided in [12] using first-order LBS approximation. In [13], a generalized formulation and first-order LBS approach was provided for a class of control-affine ES systems that: (i) unifies previous works such as [11, 12], and (ii) provides a new control law with enhanced stability guarantees. It is important to emphasize that ES methods based on first-order LBS approximation (e.g., [11, 12, 13]) are gradient-based ES approaches; that is, the ES system is designed such that its LBS/average behaves as a gradient-flow. In [14], the authors generalized further the conditions of $p_{i}$ and $k_{i}$ compared to [11], and introduced a second-order LBS, which approximates (1). This enabled the introduction of a Newton-based ES approach because the inclusion of second-order Lie brackets provided access to second-order derivative information (i.e., Hessian), hence enabling a Newton-based flow. Recently, the authors in [15] provided a generalized approach for generating LBSs with higher-orders to approximate (1). They also showed that LBS approximations are averaging terms themselves, which guarantees the closeness of trajectories between LBSs and the original system (1) given small enough $\varepsilon$. Moreover, it was observed that ES designs based on third-order LBS possess a faster convergence rate when compared with even Newton-based ES; the authors in [15] argued (without proof) that the observed faster convergence is due to the inclusion of third-order Lie brackets, which provide access to higher-order derivatives beyond Hessian (i.e., third-order derivative information).
Motivation & contributions. Inspired by the concept of higher-order Lie bracket averaging, this paper further explores the application of these techniques to extremum seeking problems, with a particular focus on achieving faster convergence rates. In particular, many extremum seeking algorithms based on first-order Lie bracket approximations exhibit exponential convergence when the cost function $J(x)$ behaves locally like a quadratic function near the extremum point $x^{*}$ (i.e., $J(x)\sim\|x-x^{*}\|^{2}$). However, if the cost function behaves like a higher-degree polynomial near the extremum, i.e., $J(x)\sim\|x-x^{*}\|^{m}$ with $m>2$, such algorithms exhibit only a polynomial decay rate. Even though it was observed in [15] that using higher-order Lie brackets to design ES may lead to faster convergence rate when $J(x)\sim\|x-x^{*}\|^{4}$, no conclusion or a proof was provided regarding the nature of the faster convergence resulting from higher-order Lie brackets.
In this paper, we investigate a class of extremum seeking problems involving the unconstrained minimization of a cost function $J(x)$. The function $J$ is unknown in terms of an explicit analytic expression, but it can be evaluated (measured) at any point. We focus on designing an ES control system of the form $\dot{x}=u(t,J(x)),$ such that the system’s trajectories tend asymptotically to an extremum point of the function $J$. To develop our approach, in this paper we assume that the cost function behaves locally like an $m$-th order polynomial near the minimizer $x^{*}\in\mathbb{R}^{n}$, i.e., $J(x)\sim\|x-x^{*}\|^{m}$, with some $m\geq 2$.
The main contribution of this paper is a novel extremum seeking design framework that leverages the excitation of higher-order Lie brackets to steer the system along directions corresponding to higher-order derivatives of the cost function. This approach helps to increase the convergence rate, in particular, ensuring the exponential convergence even for cost functions that are not quadratic in nature. Furthermore, we generalize the result of [13], which described a family of vector fields whose first-order Lie bracket equals the gradient of the cost function. In this paper, we extend this idea by deriving a formula that generates vector fields such that the corresponding $\ell$-th order Lie bracket equals the $\ell$-th derivative of the cost function. In addition, we discuss several strategies for exciting Lie brackets using different numbers of dither signals, which gives higher flexibility in the design of control vector fields.
The remainder of the paper is organized as follows. Section II introduces main notations and definitions and preliminary results related to Lie bracket-based extremum seeking control, and describes the main idea behind our approach. Section III presents the main contributions of this paper, which are illustrated through numerical simulations in Section IV. Finally, Section V provides concluding remarks and outlines potential extensions and applications of the proposed framework.
II Preliminaries
II-A Notations
Throughout the text, $\mathbb{R}^{+}=[0,\infty)$ denotes the set of all non-negative real numbers;
$B_{\delta}(x^{*})$ is the $\delta$-neighborhood of $x^{*}{\in}\mathbb{R}^{n}$, $\overline{B_{\delta}(x^{*})}$ is its closure;
for $h{\in}C^{1}(\mathbb{R}^{n};\mathbb{R})$, $\xi{\in}\mathbb{R}^{n}$, we denote $\nabla h(\xi):=\frac{\partial h(x)}{\partial x}^{T}\Big{|}_{x=\xi}$ to be a column vector;
for $h\in C^{N}(\mathbb{R};\mathbb{R})$, we define its $\ell$-th order derivative as $h^{(\ell)}(x):=\frac{d^{\ell}h(x)}{dx^{\ell}}$, for any $\ell\in\{0,\dots,N\}$, with $h^{(0)}(x):=h(x)$;
for a function $f:\mathbb{R}\to\mathbb{R}$, $f(z)=O(z)$ as $z\to 0$ means that there is a $c>0$ such that $|f(z)|\leq c|z|$ in some neighborhood of 00;
for $f,g:\mathbb{R}^{n}\to\mathbb{R}^{n}$, $x\in\mathbb{R}^{n}$, the Lie derivative is defined as $L_{g}f(x)=\lim\limits_{s\to 0}\dfrac{f(x+sg(x))-f(x)}{s}$,
and $[f,g](x)=L_{f}g(x)-L_{g}f(x)$ is the first order Lie bracket.
to define higher-order Lie brackets, we introduce $\ell$-dimensional multi-index $I_{\ell}=(i_{1},\dots,i_{\ell})$; then
$f_{I_{\ell}}(x)=\left[\big{[}[f_{i_{1}},f_{i_{2}}],f_{i_{3}}\big{]},\dots,f_{i_{\ell}}\right](x)$ – the right-iterated Lie bracket of length $\ell$, or $(\ell-1)$-order Lie bracket; for $I_{1}{\in}\{1,...,m\}$, $f_{I_{1}}$ simply means a corresponding vector field;
$L^{\infty}_{[0,\varepsilon]}$ – the class of essentially bounded measurable functions on $[0,\varepsilon]$.
II-B Lie brackets approximations
[…]
$\dot{\bar{x}}{=}f_{0}(\bar{x})+\sum_{i=1}^{r}L_{i}(\bar{x}),$ (3)
with
$L_{1}$ $=0,$
$L_{2}$ $=\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\nu_{j_{1}j_{2}}[f_{j_{1}},f_{j_{2}}],$
$L_{3}$ $=\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\sum_{j_{3}=1}^{m}\nu_{j_{1}j_{2}j_{3}}[f_{j_{3}},[f_{j_{1}},f_{j_{2}}]],$
$\begin{split}L_{4}&=\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\sum_{j_{3}=1}^{m}\sum_{j_{4}=j_{3}+1}^{m}\beta_{1_{j_{1}j_{2}j_{3}j_{4}}}\Big{[}[f_{j_{1}},f_{j_{2}}],[f_{j_{3}},f_{j_{4}}]\Big{]}\\ &+\sum_{j_{1}=1}^{m}\sum_{j_{2}=j_{1}+1}^{m}\sum_{j_{3}=1}^{m}\sum_{j_{4}=1}^{m}\beta_{2_{j_{1}j_{2}j_{3}j_{4}}}[[[f_{j_{1}},f_{j_{2}}],f_{j_{3}}],f_{j_{4}}],\end{split}$
where $\nu_{j_{1}j_{2}}$, $\nu_{j_{1}j_{2}j_{3}}$, $\beta_{1_{j_{1}j_{2}j_{3}j_{4}}}$ and $\beta_{2_{j_{1}j_{2}j_{3}j_{4}}}$ are coefficients resulting from the iterated integrals of the dither input signals (formulas are provided in [15, Section 4]). Truncating (3) at $r=2$ provides first-order LBS (e.g., [11, 13]). Similarly, truncating (3) at $r=3$ and $r=4$ provides second- and third-order LBS, respectively. The stability properties of systems (1) and (3) are related as follows.
Lemma 1 ([11, 13, 15])
If a compact set $S\subset\mathbb{R}^{n}$ is locally (globally) uniformly asymptotically stable for (3) then it is locally (semi-globally) practically uniformly asymptotically stable for (1).
Below we recall the notion of practical stability.
Definition 1
A compact set $S\subset\mathbb{R}^{n}$ is said to be locally practically uniformly asymptotically stable for (1) if:
– it is practically uniformly stable, i.e. for every $\rho>0$ there exist $\delta>0$ and $\bar{\varepsilon}>0$ such that, for all $t_{0}\geq 0$ and $\varepsilon\in(0,\bar{\varepsilon})$, any solution of (1) with $x^{0}\in B_{\delta}(S)$ satisfies $x(t)\in B_{\varepsilon}(S)$ for all $t\geq t_{0}$;
$\hat{\delta}$-practically uniformly attractive with some $\hat{\delta}>0$, i.e. for every $\rho>0$ there exist $t_{\rho}\geq 0$ and $\bar{\varepsilon}>0$ such that, for all $t_{0}\geq 0$ and $\varepsilon\in(0,\bar{\varepsilon})$, any solution of (1) with $x^{0}\in B_{\delta}(S)$ satisfies $x(t)\in B_{\varepsilon}(S)$ for all $t\geq t_{0}+t_{1}$.
If the attractivity property holds for every $\hat{\delta}{>}0$, then the set $S$ is called semi-globally practically uniformly asymptotically stable for (1).
Lemma 1 establishes the relation between the solutions of control-affine system (1) and the corresponding first-order Lie bracket system LBS (3). Let us recap the approach from [13] given its relevance to the contributions of this paper. For a special class of (1) when $f_{0}=0$, $p_{i}=0.5$, $k_{i}=1$ for all $i$, the result of Lemma 1 can be exploited for solving the extremum seeking problem in the following way [11, 13]: let us consider a class of ES systems of the form
$\dot{x}=\frac{1}{\sqrt{\varepsilon}}\Big{(}F_{1}(J(x))u_{1}^{\varepsilon}(t)+F_{2}(J(x))u_{2}^{\varepsilon}(t)\Big{)},$ (4)
where $F_{1}\circ J,F_{2}\circ J\in C^{2}(D)$ satisfy the relation $[F_{1},F_{2}](Z)=1$ for all $z\in\mathbb{R}$ (which implies $[F_{1}\circ J,F_{2}\circ J](x)=\nabla J(x)$), $u_{1}^{\varepsilon}(t)=2\sqrt{{\pi}{\varepsilon^{-1}}}\cos\big{(}{2\pi t}{\varepsilon^{-1}}\big{)}$, $u_{2}^{\varepsilon}(t)=2\sqrt{{\pi}{\varepsilon^{-1}}}\sin\big{(}{2\pi t}{\varepsilon^{-1}}\big{)}$, and assume that the cost function $J$ satisfies the following properties in some domain $D\subseteq\mathbb{R}^{n}$:
Assumption 1
The function $J\in C^{2}(D,\mathbb{R}^{n})$, and
  • A1.1)
    there exists an $x^{*}{\in}D$ such that
    $\nabla J(x)=0\text{ if and only if }x=x^{*},$
    and $J(x^{*})=J^{*}{\in}\mathbb{R}$, $J(x)>J(x^{*})$ for all $x\in D{\setminus}\{x^{*}\}$.
  • A1.2)
    There exist constants $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2},\mu$, and $m\geq 1$, such that, for all $x\in D$,
    $\alpha_{1}\|x{-}x^{*}\|^{m}\leq$ $J(x)-J^{*}\leq\alpha_{2}\|x{-}x^{*}\|^{m},$
    $\beta_{1}(J(x)-J^{*})^{1{-}\frac{1}{m}}\leq$ $\|\nabla J(x)\|\leq\beta_{2}(J(x)-J^{*})^{1{-}\frac{1}{m}},$
    $\left\|\frac{\partial^{2}J(x)}{\partial x^{2}}\right\|{\leq}$ $\mu(J(x)-J^{*})^{1{-}\frac{2}{m}}.$
Assumption A1.1) states that the cost function $J$ possesses an isolated local minimum at $x^{*}$,where it attains the value $J^{*}$. Assumption A1.2) reflects the requirement that $J$ exhibits a local behavior similar to that of a power function.
It can be shown that the point $x=x^{*}$ is practically asymptotically stable for (4), with the convergence rate dependent on the parameter $m$ in A1.2). Namely, the following result follows from [13]:
Lemma 2
If the cost function $J\in C^{2}(\mathbb{R}^{n};\mathbb{R})$ satisfies Assumption 1 in a domain $D\subset\mathbb{R}^{n}$, then $x^{*}$ is practically exponentially stable for system (4) if $m=2$, and $x^{*}$ is practically asymptotically stable for system (4) if $m>2$. Namely, for any $\delta$ such that $\overline{B_{\delta}(x^{*})}\subset D$, any $\bar{\lambda}\in(0,\alpha\kappa_{1})$, and $\rho\in(0,\delta)$, there exists an $\bar{\varepsilon}>0$ such that, for any $\varepsilon\in(0,\bar{\varepsilon}]$, $\lambda\in(0,\bar{\lambda}]$, the solutions of system (4) with $x^{0}{\in}B_{\delta}(x^{*})$ exhibit the following decay rate:
  • if $m=2$, then
    $\|x(t)-x^{*}\|\leq c_{m}\|x^{0}-x^{*}\|e^{-\lambda t}+\rho;$
  • if $m>2$, then
    $\|x(t)-x^{*}\|\leq\Big{(}c_{m}\|x^{0}-x^{*}\|^{2-m}+\tilde{c}_{m}t\Big{)}^{-1/(m-2)}+\rho.$
Here $c_{m},\tilde{c}_{m}$ are positive constants; $c_{m}$ can be made arbitrarily close to 1 by choosing a sufficiently small $\varepsilon$.
More technical details on the above decay rate estimates can be found in [13].
For the sake of clarity, we have assumed here $x\in\mathbb{R}$, however, the above result can be easily extended to the case $x\in\mathbb{R}^{n}$ (see, e.g., [11, 13]).
II-C Main idea
As follows from Lemma 2, the Lie bracket approximation approach provides a constructive solution to the extremum seeking problem which ensures the exponential convergence of the trajectories of system (4) to the optimal point in the case of a quadratic-like cost function, i.e., for $m=2$ in Assumption A2). However, for cases $m>2$, the above algorithm ensures only a polynomial decay rate. This can also be observed by analyzing the first-order Lie bracket system associated with system (4):
$\bar{x}=-\nabla J(\bar{x}).$
Note that the above first-order LBS is a complete average asymptote for (4), meaning that higher-order LBS approximations will be redundant [15]. Now, if for example $J=\dfrac{1}{2}(x-x^{*})^{2}$, then the Lie bracket system is linear, $\dot{\bar{x}}=-(\bar{x}-x^{*})$, and thus $x^{*}$ is its exponentially stable equilibrium point. In case $J=\dfrac{1}{4}(x-x^{*})^{4}$, the Lie bracket system takes the form $\dot{\bar{x}}=-(\bar{x}-x^{*})^{3}$, and its solutions are well-known to exhibit the polynomial decay rate $O(t^{-1/2})$ as $t\to\infty$ [30, 31]. Assume now that we can associate the properties of system (1) with a system which gives access to the third-order derivative of $J$, namely, with the system $\dot{\tilde{x}}=-J^{(3)}(\tilde{x})=-6(\tilde{x}-x^{*})$. Then the latter system turns out to be linear again, which, under certain assumptions, may imply the practical exponential stability of the extremum seeking system. A natural way of accessing higher order derivatives of the cost function is to excite the Lie brackets of corresponding order [14, 15]. For example, it is easy to see that
$[[[J,1],1],1](x)=-J^{(3)}(x).$
The main idea of this paper is to construct an extremum seeking system in the form
$\dot{x}=\sum_{k=1}^{n_{u}}g_{k}(J(x))u_{k}^{\varepsilon}(t),$ (5)
so that, under a special choice of control vector fields $g_{k}$, $k=1,\dots,n_{u}\in\mathbb{N}$, dither signals $u_{k}^{\varepsilon}\in L^{\infty}_{[0,\varepsilon]}$, and a parameter $\varepsilon>0$, its trajectories approximate the trajectories of a system with high order Lie brackets:
$\dot{\bar{x}}=\sum_{\ell=1}^{N}\sum_{I_{\ell}\in S_{\ell}}c_{I_{\ell}}g_{I_{\ell}}(\bar{x}),$ (6)
where $N\in\mathbb{N}$, $S_{\ell}\subset\{1,\dots,n_{u}\}^{\ell}$ denotes the sets of of multi-indices of the Lie brackets required for solving the extremum seeking problem, $g_{I_{\ell}}$ are the corresponding Lie brackets, and $c_{I_{\ell}}$ are constant parameters. For example, for system (3) with $r=2$, $f_{0}=0$, $f_{i}=g_{i}$, $i\in\{1,\dots,n\}$, we mean $N=2$, $S_{1}=\emptyset$, $S_{2}=\{(i,j):1\leq i<j\leq n\}$, $I_{2}:=(i,j)\in S_{2}$ $c_{I_{2}}=\nu_{ij}$.
One of the main tools exploited in this paper is the Chen–Fliess series expansion: under certain regularity assumptions on the control vector fields of system (5), the solutions of system (5) with $x(0)=x^{0}$ can be represented in the following form [32, 19]:
$x(t)=x^{0}+$ $\sum_{\ell=1}^{N}\sum_{k_{1},\dots,k_{\ell}=1}^{n_{u}}L_{g_{k_{\ell}}}\dots L_{g_{k_{2}}}f_{g_{1}}(x^{0})$ (7)
×$\times\int\limits_{0}^{t}\int\limits_{0}^{s_{1}}...\int\limits_{0}^{s_{\ell-1}}u_{k_{1}}^{\varepsilon}(s_{1})\times\dots$
×$\qquad\qquad\qquad\times u_{k_{\ell}}^{\varepsilon}(s_{\ell})ds_{\ell}...ds_{1}+R(t),$
with the remainder
$R(t)=$ $\sum_{k_{1},\dots,k_{N+1}=1}^{n_{u}}\int\limits_{0}^{t}\int\limits_{0}^{s_{1}}...\int\limits_{0}^{s_{N}}L_{g_{k_{N+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}(x(s_{N+1}))$
×$\times u_{k_{1}}^{\varepsilon}(s_{1})\times\dots\times u_{k_{N+1}}^{\varepsilon}(s_{N+1})ds_{N+1}...ds_{1}.$
III Main results
III-A Two-input extremum seeking system
To simplify the presentation, in this section, we assume $x\in\mathbb{R}$. To steer the solutions of an extremum seeking system towards the direction of high-order Lie brackets, we refer to control approaches from nonholonomic systems theory [9, 10, 33, 17, 34, 35, 19, 20]. We focus here on the two-input systems of form (5):
$\dot{x}=g_{1}(J(x))u_{1}^{\varepsilon}(t)+g_{2}(J(x))u_{2}^{\varepsilon}(t).$ (8)
Suppose also that $u_{1}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{1}(t/\varepsilon)$, $u_{2}^{\varepsilon}(t)={\varepsilon^{1/N-1}}v_{2}(t/\varepsilon)$, with the dithers $v_{1}(t/\varepsilon),v_{2}(t/\varepsilon)$ exciting the Lie bracket
$g_{I_{N}}(z)=\big{[}\big{[}\dots[g_{1},\underbrace{g_{2}],g_{2}\big{]},\dots,g_{2}}_{N-1\text{ times}}]\dots]\big{]}(z)$
at time $t=\varepsilon$, with $I_{N}=(1,\underbrace{2,\dots,2}_{N-1\text{ times}})$, in the sense that the Chen–Fliess series expansion (7) takes the form
$x(\varepsilon)=x^{0}+\varepsilon g_{I_{N}}(J(x^{0}))+R(\varepsilon),$ (9)
and all the other Lie brackets of length from 1 to $N$ do not appear in the above expansion. One way to construct such inputs is described in [34, 35], other approaches can be found in, e.g., [9, 10, 15]. The concrete examples are as follows, as mentioned in [20].
Statement 1
  • the inputs $v_{1}(t/\varepsilon)=v_{1}^{12}(t/\varepsilon)=2\sqrt{\kappa_{12}\pi}\cos\left({2\kappa_{12}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=v_{2}^{12}(t/\varepsilon)=2\sqrt{{\kappa_{12}\pi}}\sin\left({2\kappa_{12}\pi t}/{\varepsilon}\right)$, $\kappa_{12}\in\mathbb{Z}$,
    excite the first order Lie bracket $[g_{1},g_{2}]$;
  • the inputs
    $v_{1}(t/\varepsilon)=v_{1}^{122}(t/\varepsilon)=-2\left({4\kappa_{122}\pi}\right)^{\frac{2}{3}}\cos\left({4\kappa_{122}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=v_{2}^{122}(t/\varepsilon)=\left({4\kappa_{122}\pi}\right)^{\frac{2}{3}}|\cos\left({2\kappa_{122}\pi t}/{\varepsilon}\right)$,
    $\kappa_{112}\in\mathbb{Z}$, excite the second order Lie bracket $[[g_{1},g_{2}],g_{2}]$;
  • the inputs
    $v_{1}(t/\varepsilon)=v_{1}^{1222}(t/\varepsilon)=6\left(2\kappa_{1222}\pi\right)^{\frac{3}{4}}\sin\left({6\kappa_{1222}\pi t}/{\varepsilon}\right)$,
    $v_{2}(t/\varepsilon)=v_{2}^{1222}(t/\varepsilon)=2\left(2\kappa_{1222}\pi\right)^{\frac{3}{4}}\cos\left({2\kappa_{1222}\pi t}/{\varepsilon}\right)$,
    $\kappa_{1222}\in\mathbb{Z}$, excite the third order Lie bracket $\big{[}[[g_{1},g_{2}],g_{2}],g_{2}\big{]}$.
Remark 1
By examining (3) in light of the design choices in Statement 1 above, we note:
  • the choice of $v_{1}^{12}$ and $v_{2}^{12}$ to excite $[g_{1},g_{2}]$ fits the condition provided in [15, Section 4.1] when $p_{1}+p_{2}=1/2+1/2=1$, which guarantees bounded first-order LBS as $\varepsilon\rightarrow 0$ and that the realized first-order LBS is a complete average asymptote (i.e., higher-order LBS approximations are made redundant by that design in that they vanish as $\varepsilon\rightarrow 0$).
  • the choices of $v_{1}^{122}$ and $v_{2}^{122}$ to excite $[[g_{1},g_{2}],g_{2}]$ fits the condition provided in [15, Section 4.1] when $p_{1}+p_{2}+p_{2}=2/3+2/3+2/3=2$, which means that the second-order LBS is designed as the complete average asymptote. Also, due to the design choice of the dither signals with $k_{1}=4\kappa_{122}$ and $k_{2}=2\kappa_{122}$, coefficients resulting from the iterated integrals for first-order terms and other second-order terms (except the excited bracket) vanish. This also leads to a bounded second-order LBS as $\varepsilon\rightarrow 0$.
  • the choices of $v_{1}^{1222}$ and $v_{2}^{1222}$ to excite $\big{[}[[g_{1},g_{2}],g_{2}],g_{2}\big{]}$ fits the condition provided in [15, Section 4.1] when $p_{1}+p_{2}+p_{2}+p_{2}=3/4+3/4+3/4+3/4=3$, which means that the third-order LBS is designed as the complete average asymptote. Also, due to the design choice of the dither signals with $k_{1}=6\kappa_{1222}$ and $k_{2}=2\kappa_{1222}$, coefficients resulting from the iterated integrals for first-order terms, second-order terms and other third-order terms (except the excited bracket) vanish. This also leads to a bounded third-order LBS as $\varepsilon\rightarrow 0$.
Assume further that $g_{1},g_{2}$ are chosen in such a way that the Lie bracket $g_{I_{N}}(J(x))$ has the form
$g_{I_{N}}(J(x))=-c_{N}J^{(N-1)}(x),$ (10)
and some $c_{N}>0$ playing a role of control gain parameter. In this context, $g_{I_{N}}(J(x))$ denotes the corresponding Lie bracket computed with the compositions of functions $g_{1}\circ J$, $g_{2}\circ J$. The most obvious choice of the vector fields satisfying relation (10) is $g_{1}(z)=(-1)^{N+1}z$, $g_{2}(z)\equiv 1$. For the first-order Lie brackets, the whole family of functions $g_{1},g_{2}$ satisfying the relation
$[g_{1}\circ J,g_{2}\circ J](x)=-\varphi(J(x))\nabla J(x),$ (11)
with any given continuous function $\varphi$, has been introduced in [13]. Then the expansion (9) takes the form
$x(\varepsilon)=x^{0}-\varepsilon c_{N}J^{(N-1)}(x^{0})+R(\varepsilon),$ (12)
and, similarly to the approach of [13], the practical asymptotic stability of $x^{*}$ for system (8) can be proved. We proceed by summarizing the key results of this subsection and integrating them into the context of solving the extremum seeking problem. For this purpose, we further specify the properties of $J$ as a polynomial-like single-variable function:
Assumption 2
The function $J\in C^{m}(D,\mathbb{R}^{m})$ with some $m\geq 2$, and there exist constants $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$, such that, for all $x\in D$,
$\alpha_{1}\|x-x^{*}\|^{m}\leq J(x)-J^{*}$ $\leq\alpha_{2}\|x-x^{*}\|^{m},$
$J^{(m-1)}(x)(x-x^{*})$ $\geq\beta_{1}\|x^{0}-x^{*}\|^{2},$
$\|J^{(m-1)}(x)\|$ $\leq\beta_{2}\|x^{0}-x^{*}\|.$
Theorem 1
Given system (8) with $N=m$ and a cost function $J$ satisfying Assumption 2, let the vector fields $g_{1}\circ J,g_{2}\circ J\in C^{m}(D;\mathbb{R}^{n})$ satisfy the relation (10) with $I_{m}=(1,{2,\dots,2})$, and let $u_{1}^{\varepsilon},u_{2}^{\varepsilon}\in L^{\infty}_{[0,\varepsilon]}$ be $\varepsilon$-periodic dithers which ensure the representation (9). Then the point $x^{*}$ is semi-globally practically exponentially stable for system (8).
Proof:
The argumentation is similar to the proof of the practical exponential stability in [13, Theorem 3], so we only explain here how to derive exponential decay rate to a neighborhood of $x^{*}$ based on the representation (9). Given any $\delta>0$, let $D^{\prime}$ be any compact set such that $\overline{B_{\delta}(x^{*})}\subset D^{\prime}\subset D$. Similarly to the proof ofn [13, Theorem 3], one can show that there exists an $\varepsilon_{0}>0$ such that solutions of system (8) with $x^{0}\in B_{\delta}(x^{*})$ and $\varepsilon\in(0,\varepsilon_{0})$ are well-defined in $D^{\prime}$ for all $t\in[0,\varepsilon]$. To estimate the remainder in (12), denote
$c_{Fm}=\sup\limits_{x\in D^{\prime}}\sum_{k_{1},\dots,k_{m+1}=1}^{2}L_{g_{k_{m}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J(x),$
$c_{u}=\max\limits_{k=1,2}\sup\limits_{t\in[0,\varepsilon]}|v_{k}(t/\varepsilon)|$. Then it is easy to see that
$\|R(\varepsilon)\|=$
$\sum_{k_{1},\dots,k_{m+1}=1}^{2}$ $\int\limits_{0}^{t}\int\limits_{0}^{s_{1}}...\int\limits_{0}^{s_{m}}L_{g_{k_{m+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J(x(s_{m+1}))$
×$\times u_{k_{1}}^{\varepsilon}(s_{1})\times\dots\times u_{k_{m+1}}^{\varepsilon}(s_{m})ds_{m}...ds_{1}$
$\leq$ $c_{R}\varepsilon^{1+\frac{1}{m}},$
where $c_{R}=c_{u}^{m+1}c_{Fm}$. The above estimate together with the representation (12), Assumption 2, and triangular inequality, implies that
$\|x(\varepsilon)-x^{*}\|^{2}\leq$ $\|x^{0}-x^{*}\|^{2}(1-2\varepsilon\gamma_{m})+\varepsilon^{1+\frac{1}{m}}\sigma,$
with $\gamma_{m}=c_{m}(\beta_{1}-\varepsilon c_{m}\beta_{2}^{2})$, $\sigma=c_{R}^{2}\varepsilon^{1+\frac{1}{m}}+2c_{R}\delta(1+\varepsilon c_{m}\beta_{2})$. Then following the argumentation of [13, Steps 3.I-4.I], we can establish the practical exponential decay rate to an arbitrary small neighborhood of $x^{*}$. ∎
Remark 2
As in the paper [13], it is also possible to relax regularity assumption on the control vector fields. Namely, requirement of $g_{k}\circ J$, $k=1,2$, being $m$ times continuously differentiable in $D$ can be replaced with the following: $g_{k}\circ J\in C^{m}(D\setminus\{x^{*}\});\mathbb{R}$ and $L_{g_{k_{N+1}}}\dots L_{g_{k_{2}}}g_{k_{1}}\circ J\in C(D;\mathbb{R})$ for all $N\in\{1,\dots,m\}$, $k_{1},\dots,k_{m+1}\in\{1,2\}$. This relaxation is particularly important for deriving conditions for the “classical” exponential stability in the sense of Lyapunov, meaning that the trajectories of the extremum seeking system converge to the point $x^{*}$, rather than merely to its neighborhood. Another important condition for achieving classical exponential stability is the property of vanishing amplitudes, which requires that $g_{k}\circ J\to 0$ as $x\to x^{*}$ (see [13, Theorem 3, Part II]). Since selecting such vector fields becomes increasingly challenging in the case of higher-order Lie brackets, we leave this task, along with a rigorous formulation of the corresponding exponential stability properties, for furture research.
III-B Alternative design approaches
In the previous subsection, we have described the approach for generating extremum seeking systems with two inputs (dithers). However, this method for realizing the dynamics similar to $\dot{\bar{x}}=-J^{(m-1)}$ is clearly not unique. For example, an alternative way to excite a Lie bracket of length 2 is to introduce a three-input strategy
$\dot{x}=g_{1}(J(x))u_{1}^{\varepsilon}(t)+g_{2}(J(x))u_{2}^{\varepsilon}(t)+g_{3}(J(x))u_{3}^{\varepsilon}(t),$ (13)
with $u_{k}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{k}^{\varepsilon}(t)$, $k=1,2,3$, with $v_{k}(t/\varepsilon)$ exciting the Lie bracket $g_{(1,2,3)}=[[g_{1},g_{2}],g_{3}]$, and $g_{1},g_{2},g_{3}$ such that
$\big{[}[g_{1}\circ J,g_{2}\circ J],g_{3}\circ J\big{]}(x)=-c_{(1,2,3)}J^{(2)}(x)$ (14)
with some $c_{(1,2,3)}>0$. The reason for introducing more inputs in an extremum seeking system is to gain more flexibility in selecting the control vector fields. As in previous subsection, formula (14) holds for $g_{1}(z)=-z$, $g_{2}(z)=1$, $g_{3}(z)=1$. The three-input structure, however, facilitates a more general description of the entire class of functions $g_{1},g_{2},g_{3}$ satisfying (14). Namely, the paper [13] provides a general formula for deriving $g_{1},g_{2}$ such that the relation (11) is satisfied with an arbitrary $\varphi:\mathbb{R}\to\mathbb{R}$. Then straightforward computations yields the following result.
Lemma 3
Let $\varphi_{2}\in C^{1}(\mathbb{R};\mathbb{R})$ be any given function, the functions $g_{1},g_{2}\in C^{1}(\mathbb{R};\mathbb{R})$ satisfy (11) with $\varphi=\varphi_{2}$, and let $g_{3}=-\varphi_{2}$. Then
$\big{[}[g_{1}\circ J,g_{2}\circ J],g_{3}\circ J\big{]}(x)=-\varphi_{2}^{2}(x)J^{(2)}(x)\text{ for all }x\in D.$
Proof:
Indeed, under the above conditions,
[…]
$=\big{[}\varphi_{2}(x)\nabla J(x),\varphi_{2}(J(x))\big{]}(x)=-\varphi_{2}^{2}(J(x))J^{(2)}(x).$
In particular, by setting $\varphi_{2}(z)\equiv\sqrt{c_{(1,2,3)}}$ in the above Lemma, we directly obtain formula (14).
As noted in Remark 1, the regularity assumptions on $g_{1},g_{2},g_{3}$ can be relaxed, and with appropriate selections, it is possible to achieve exponential stability in the sense of Lyapunov. A formal statement of this result is left for future work.
Remark 3
With the exception of special cases (e.g., Newton-based ES as in [14]), in general, the even-order derivatives are rather not helpful for classic extremum seeking problems. The goal of considering this case here is to illustrate alternative approaches to designing extremum seeking systems, both in terms of selecting control vector fields and determining the number of dithers. The formula presented in Lemma 3 can be easily extended to higher-order scenarios. For example, given a four inputs system (5) with $n_{u}=4$ and any function $\varphi_{3}\in C^{1}(\mathbb{R};\mathbb{R}^{+})$ such that $\sqrt{\varphi_{3}}\in C^{1}(\mathbb{R}^{+};\mathbb{R}^{+})$, let the functions $g_{1},g_{2},g_{3}$ be chosen as in Lemma 3 with $\varphi_{2}=\sqrt{\varphi_{3}}$. Then $\big{[}\big{[}[g_{1}\circ J,g_{2}\circ J],g_{3}\circ J\big{]},g_{4}\circ J\big{]}(x)=-\varphi_{3}^{2}J^{(3)}(x)$.
To conclude this subsection, we provide a generalized formulation of Theorem 1.
Theorem 2
Given system (8) with $N=m$ and a cost function $J$ satisfying Assumption 2, let the vector fields $g_{1}\circ J,\dots,g_{m}\circ J\in C^{m}(D;\mathbb{R}^{n})$ satisfy the relation (10) with $I_{m}=(i_{1},\dots,i_{m})\in\{1,\dots,m\}^{m}$, and let $u_{1}^{\varepsilon},\dots,u_{m}^{\varepsilon}\in L^{\infty}_{[0,\varepsilon]}$ be $\varepsilon$-periodic dithers which ensure the representation (9). Then the point $x^{*}$ is semi-globally practically exponentially stable for system (8).
The proof mostly repeats the proof of Theorem 1.
III-C Polynomial-like cost functions with unknown degree
Ensuring dynamics of the form $\dot{\bar{x}}=-J^{(N-1)}$ is particularly useful when the degree $m$ characterizing the behavior of the cost function is known. However, in extremum seeking problems, only very limited information about the cost function is typically available. In such cases, an incorrect choice of the order $N$ may lead to undesirable system behavior, making it impossible to steer system to the desired state. For example, consider the cost function $J(x)=\frac{1}{2}(x-x^{*})^{2}$, and assume that inputs exciting the third-order Lie bracket are used in (8). The resulting third-order Lie bracket system then takes the form $\dot{\bar{x}}=0$, which loses the desired stability properties. Correspondingly, the expansion (12) becomes $x(\varepsilon)=x^{0}+R(\varepsilon),$ which also demonstrates the ineffectiveness of such control in solving the extremum seeking problem in this case.
To overcome this limitation, we employ a splitting of the time-varying dithers, where different dithers are assigned to excite Lie brackets of different orders. A clever choice of dither’s frequencies allows to obtain an associated Lie bracket system of the form
$\dot{\bar{x}}=-\sum_{j=1}^{N}\gamma_{j}J^{(j)}(\bar{x})\text{ with some }\gamma_{j}\geq 0,$
which, under appropriate assumptions on $J$, possesses local exponential stability in a neighborhood of $x^{*}$. To illustrate the method clearly and stay within page limits, we excite each Lie bracket with two corresponding inputs, and restrict our study to $N=3$.
Assumption 3
The function $J\in C^{m}(D,\mathbb{R}^{m})$ with some $m\geq 2$, and there exist constants $\alpha_{1},\alpha_{2}>0$, $\beta_{\ell 1},\beta_{\ell 2}\geq 0$, $\ell=1,2$, such that, for all $x\in D$,
$\alpha_{m1}\|x-x^{*}\|^{m}\leq J(x)-J^{*}$ $\leq\alpha_{m2}\|x-x^{*}\|^{m},$
$J^{(1)}(x)(x-x^{*})$ $\geq\beta_{11}\|x^{0}-x^{*}\|^{m},$
$\|J^{(1)}(x)\|$ $\leq\beta_{12}\|x^{0}-x^{*}\|^{m-1},$
$J^{(3)}(x)(x-x^{*})$ $\geq\beta_{21}\|x^{0}-x^{*}\|^{m-2},$
$\|J^{(3)}(x)\|$ $\leq\beta_{22},$
assuming $\beta_{21}=0$ if $m=2$ and $\beta_{21}\geq 0$ otherwise.
Theorem 3
Given a cost function satisfying Assumption 3, let the extremum seeking system have the form
$\dot{x}=\sum_{k=1}^{4}g_{k}(J(x))u_{k}^{\varepsilon}(t),$ (15)
where
$u_{1}^{\varepsilon}(t)=\varepsilon^{-1/2}v_{1}^{12}(t/\varepsilon),\,u_{2}^{\varepsilon}(t)=\varepsilon^{-1/2}v_{2}^{12}(t/\varepsilon),$
$u_{3}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{1}^{1222}(t/\varepsilon),\,u_{4}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{2}^{1222}(t/\varepsilon),$
with the dithers $v_{j}^{12},v_{j}^{1222}$, $j=1,2$, chosen as in Statement 1 under assumption that there are no resonances of order up to $4$ in each of the following pairs: $(\kappa_{12},\kappa_{1222})$, $(\kappa_{12},3\kappa_{1222})$, and $(\kappa_{12},3\kappa_{1222})$.
Further, assume that the functions $g_{k}\circ J\in C^{4}(D;\mathbb{R})$ satisfy the following relations:
$[g_{1}\circ J,g_{2}\circ J](x)=-\gamma_{1}J^{(1)}(x),$
$\big{[}\big{[}[g_{3}\circ J,g_{4}\circ J],g_{4}\circ J\big{]},g_{4}\circ J\big{]}(x)=-\gamma_{3}J^{(3)}(x).$
Then the point $x^{*}$ is practically exponentially stable for system (15) if $m=2$ or $m=4$, and practically asymptotically stable for system (15) if $m>4$.
We omit the proof because of the space limits. It is similar to the proof of Theorem 1 and uses the fact that, as it is shown in [20], the non-resonance assumption implies that the Chen–Fliess expansion (7) for system (15) takes the form
$x(\varepsilon)=x^{0}-\gamma_{1}J^{(1)}(x^{0})-\gamma_{2}J^{(3)}(x^{0})+R(\varepsilon),$
with $\|R(\varepsilon)\|\leq C_{R}\varepsilon^{5/4}$ with some $C_{R}\geq 0$, and thus,
[…]
$-2\varepsilon(\gamma_{1}J^{(1)}(x^{0})+\gamma_{2}J^{(2)}(x^{0})))(x^{0}-x^{*})+\tilde{R}(\varepsilon)$
$\leq\|x^{0}-x^{*}\|^{2}-2\varepsilon(\beta_{11}\|x^{0}-x^{*}\|^{m}$
$\qquad+\beta_{21}\|x^{0}-x^{*}\|^{m-2})+\tilde{R}(\varepsilon),$
where $\tilde{R}(\varepsilon)=o(\varepsilon^{5/4})$ as $\varepsilon\to 0$.
Figure 1: Blue: the proposed ES system in (17), which converges practically in an exponential rate to 1, the minimum point of $J_{1}(x)=(x-1)^{4}$; green: a traditional ES approach from literature (16), which converges practically to 1 in a polynomial rate.
IV Numerical simulations
To demonstrate the effectiveness of the proposed approach, we take a cost function as the fourth-order polynomial,
$J(x)=(x-1)^{4},$
and apply first-order-based approach [11],
$\dot{x}=2\sqrt{\frac{\pi}{\varepsilon}}\left(J_{1}(x)\cos\frac{2\pi t}{\varepsilon}+\sin\frac{2\pi t}{\varepsilon}\right),$ (16)
and fourth-order-based approach with two inputs defined as in Statement 1:
$\dot{x}=2\left(\frac{2\pi}{\varepsilon}\right)^{3/4}\left(3J_{1}(x)\sin\frac{6\pi t}{\varepsilon}+\cos\frac{2\pi t}{\varepsilon}\right).$ (17)
We initiate both equations at $x(0)=0$ and put $\varepsilon=10^{-4}$. The results of the numerical simulations are shown in Fig.1, demonstrating a significantly faster convergence rate for the solutions of system (17) compared to (16), while also exhibiting a higher amplitude of oscillations. In our future studies, we plan to explore alternative choices for the functions $g_{1}$ and $g_{2}$, specially, with vanishing amplitudes similarly to as proposed in [13], which have been shown to improve the qualitative behavior of the solutions. Another promising approach for reducing oscillations involves the use of time-varying gain techniques, such as in [36, 37, 5].
V Conclusions & Future Work
In this paper, we have introduced a novel framework for extremum seeking control that utilizes higher-order Lie bracket approximations to achieve improved convergence properties, particularly for cost functions with polynomial-like behavior near their minima. Unlike many ES approaches that rely on first-order Lie bracket systems and yield exponential convergence only for quadratic-like cost functions, the proposed approach may ensure the exponential convergence even when the cost function behaves locally like $\|x-x^{*}\|^{m}$ with $m>2$. This is achieved by integrating ideas from differential geometric control theory related to higher-order Lie bracket averaging for control-affine systems [15] with control design techniques for high-order nonholonomic systems [20, 34], and advanced analytical tools for studying dynamical systems, in particular, the Chen–Fliess series expansion [13, 19].
Let us note that, to better introduce our approach and due to space limitations, we have considered only the case of a single-variable cost function. In our future work, we plan to extend this approach to multi-variable cost functions under less restrictive assumptions on their local behavior, as well as to general extremum seeking problem statement like in [38, 39]. Another important research direction concerns exploring the possible choices of generating vector fields in Lemma 3. We expect that, as in [13], it is possible to achieve asymptotic (in particular, exponential) stability in the sense of Lyapunov by appropriately choosing vector fields that vanish at the extremum. In such scenarios, an improved convergence rate would be even more beneficial, as the amplitude of oscillations would also vanish.
All in all, we believe that this paper opens new possibilities for designing extremum seeking control laws with enhanced flexibility, and initiates a promising direction for further developments of high-order Lie bracket methods in optimization tasks.